# REVISOR: Beyond Textual Reflection, Towards Multimodal Introspective Reasoning in Long-Form Video Understanding

### v1

@@ 2.1 Dilemma of Text-Based Reflection
As shown in the middle part of Fig. 2, applying these text-based reflection mechanisms to long-form video understanding tasks paradoxically leads to a decline in model performance. To eliminate the possibility that this degradation stems from VL-Rethinker’s lack of video training, we developed a comparable model equipped with a purely text-based reflection mechanism and trained it on video data. Our experiments revealed that this model also failed to improve performance in long-form video understanding tasks.
@@ 3.2 Dual Attribution Decoupled Reward
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the balance of the rewards. The term $R_{final}$ is a standard task-specific reward that validates the correctness of the final refined answer $A_{final}$ against the ground truth answer $A^{*}$.
@@ 4 Experiments
The performance of the REVISOR framework for long-form video understanding tasks is then presented in Sec. 4.2. Subsequently, Sec. 4.3 conducts an ablation study to evaluate the contribution of REVISOR’s various components. Further experimental setup details, additional ablation studies.
@@ 4.1 Experimental Setup
Dataset Construction. We use a total of 25K training samples collected from STAR [li2025star], PerceptionTest [patraucean2023perception], NExT-QA [xiao2021next], CLEVRER [yi2019clevrer], LLaVA-Video-178K [zhang2024video], TimeRFT [wang2025time], CG-Bench [chen2024cg], and ReXTime [chen2024rextime].
Hyperparameter Selection. In our experiments, we set $\lambda_{1}=0.6$ and $\lambda_{2}=0.3$. We adopt AdamW [loshchilov2017decoupled] as the optimizer with a learning rate of $1\times 10^{-6}$ and a batch size of $32$. The number of rollouts is set to $8$. During both training and evaluation, the total number of video tokens in the input is limited to a maximum of $8192$.
@@ 4.2 Main Results
4.3 Ablation Study
Dual Attribution Decoupled Reward. We perform an in-depth investigation of the DADR mechanism. As shown in Tab. 3, setting $\lambda_{2}$ to 0 (i.e., using only the Final Answer Verification Reward) results in a drop of REVISOR’s Video-MME score from 65.7% to 62.2%, which is below the 64.3% achieved by the base model Qwen2.5-VL-7B. This result suggests that, without the CSSR component, the model fails to learn how to locate the correct review segment $S$ from sparse reward signals.
@@ 6 Related Work
Long-form Video Understanding. In long-form video understanding, frame selection is crucial for reducing redundancy and computational load. Traditional approaches relied on rule-based heuristics for frame selection, but increasing complexity of multimodal QA demands more adaptive strategies. Existing methods fall into three categories: external-model augmentation, agentic methods, and model-internalized selection. External-model augmentation techniques [man2025adacm2understandingextremelylongterm, shen2024longvu], such as those in VideoChat-Flash [li2025videochatflashhierarchicalcompressionlongcontext] and KeyVideoLLM [liang2024keyvideollmlargescalevideokeyframe], employ vision-text similarity models like CLIP for token compression or keyframe selection. Although efficient, these methods depend on static, one-time selection, limiting adaptability to complex queries. Agentic approaches [wang2025videochata1thinkinglongvideos, liu2025videomindchainofloraagentlong] tackle this via iterative frameworks that combine LLMs, VLMs, and CLIP to dynamically refine frame selection. For instance, VideoAgent [wang2024videoagentlongformvideounderstanding] uses an LLM as a central agent in a state-action-observation loop, while VideoTree [wang2025videotreeadaptivetreebasedvideo] employs a hierarchical tree for coarse-to-fine search, focusing computation on relevant segments. More recent model-internalized selection methods [tong2025thinkingvideovideogeneration] integrate frame selection into the model’s reasoning process. GenS [yao2025generativeframesamplerlong] trains a lightweight generative sampler for end-to-end query-relevant frame identification, yet keeps selection and QA separate. In contrast, our method reviews video segments relevant to the question during the inference process, offering strong support for the reasoning mechanism and significantly improving the model’s reasoning ability in long-form video understanding by directly integrating visual information into the reflection process.
Self-Reflection mechanism in Large Language Model. Incorporating reflection into inference has been shown to markedly improve model performance. In language models, DeepSeek-R1[guo2025deepseek] exhibits emergent self-reflective behavior that improves answer performance; Self-Refine[madaan2023self] introduces an iterative self-refinement loop in which models generate feedback and revise their outputs without extra training or external supervision; SCoRe[kumar2024training] leverages multi-round reinforcement learning to imbue reflective capabilities, delivering substantial performance improvements; GEPA[agrawal2025gepa] proposes a reflective prompt-evolution strategy with Pareto-based selection that outperforms reinforcement learning in both sample efficiency and final task performance in compound AI settings. Building on these successes, reflection has become a focal point for vision–language reasoning. VL-Rethinker[wang2025vl] imposes a pre-answer self-verification-and-correction stage via reinforcement learning, markedly improving multimodal reasoning. Look-Back[yang2025look] encourages multimodal LLMs to autonomously revisit visual inputs during reasoning, using a simple format-based reward to induce late-stage refocusing on the image without explicit visual re-injection, thereby strengthening visual fusion and perceptual grounding. Reflection-V[jian2025look] designs visual reflection patterns and introduces a visual-attention reward to sustain attention and promote repeated use of visual evidence throughout generation, substantially boosting multimodal reasoning. SRPO[wan2025srpo] further enhances self-reflection and self-correction by curating reflection-oriented supervised fine-tuning data and employing a reflection-aware reward under group relative policy optimization, significantly improving complex reasoning performance. Taken together, these lines of work underscore reflection—whether emergent or explicitly enforced via reward shaping, prompt evolution, or RL—as a general principle for advancing both language and vision–language models.

### v2

@@ 1 Introduction
To address these challenges, we propose the concept of multimodal self-reflection, which extends traditional text-based reflection into an integrated process encompassing both language and vision. This concept envisions a model that not only reviews and refines its textual reasoning but also revisits key video segments during reflection. In doing so, the model achieves joint correction of both its linguistic reasoning logic and its visual attention focus. As demonstrated in Sec. 2.2, incorporating ground-truth key video segments during reflection significantly improves accuracy on long-form video understanding tasks, highlighting the effectiveness and promise of multimodal self-reflection for enhanced reasoning in long-form video comprehension.
@@ 2.1 Dilemma of Text-Based Reflection
As shown in the middle part of Fig. 2, applying these text-based reflection mechanisms to long-form video understanding tasks paradoxically leads to a decline in model performance. To eliminate the possibility that this degradation stems from VL-Rethinker’s lack of video training, we developed a comparable model equipped with a purely text-based reflection mechanism and trained it on video data. Details are provided in Appendix Sec. 9. Our experiments revealed that this model also failed to improve performance in long-form video understanding tasks.
@@ 3.2 Dual Attribution Decoupled Reward
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the balance of the rewards.
@@
Final Answer Verification Reward. The term $R_{final}$ is a standard task-specific reward that validates the correctness of the final refined answer $A_{final}$ against the ground truth answer $A^{*}$. This reward directly targets the accuracy objective of the reflective stage (Stage 2):
$R_{final}=\mathbb{I}(A_{final}=A^{*}),$ (6)
where $\mathbb{I}(\cdot)$ is the indicator function.
@@ 4 Experiments
We first detail our experimental setup in Sec. 8.1. The performance of the REVISOR framework for long-form video understanding tasks is then presented in Sec. 4.2. Subsequently, Sec. 4.3 conducts an ablation study to evaluate the contribution of REVISOR’s various components. Please refer to Appendix Sec. 10 for the Case Study of REVISOR.
@@ 4.1 Experimental Setup
Dataset Construction. We use a total of 25K training samples collected from STAR [li2025star], PerceptionTest [patraucean2023perception], NExT-QA [xiao2021next], CLEVRER [yi2019clevrer], LLaVA-Video-178K [zhang2024video], TimeRFT [wang2025time], CG-Bench [chen2024cg], and ReXTime [chen2024rextime]. Detailed information regarding the dataset selection procedure is available in Appendix Sec. 8.2.
Hyperparameter Selection. In our experiments, we set $\lambda_{1}=0.6$ and $\lambda_{2}=0.3$. We adopt AdamW [loshchilov2017decoupled] as the optimizer with a learning rate of $1\times 10^{-6}$ and a batch size of $32$. The number of rollouts is set to $8$. During both training and evaluation, the total number of video tokens in the input is limited to a maximum of $8192$. For more details, please refer to Appendix Sec. 8.1
@@ 4.2 Main Results
4.3 Ablation Study
Dual Attribution Decoupled Reward. We perform an in-depth investigation of the DADR mechanism. As shown in Tab. 3, setting $\lambda_{2}$ to 0 (i.e., using only the Final Answer Verification Reward) results in a drop of REVISOR’s Video-MME score from 65.7% to 62.2%, which is below the 64.3% achieved by the base model Qwen2.5-VL-7B. This result suggests that, without the CSSR component, the model fails to learn how to locate the correct review segment $S$ from sparse reward signals. Please refer to Appendix Sec. 8.3 for more ablation study results.
@@ 6 Related Work
Long-Form Video Understanding. In long-form video understanding, the visual inputs are far more complex than those in image-based tasks. Consequently, identifying and distilling only the question-relevant information from large amounts of redundant visual content becomes crucial for improving model performance. Existing approaches can be grouped into three categories: external-model augmentation, agentic methods, and model-internalized selection.
@@
External-model augmentation techniques [man2025adacm2understandingextremelylongterm, shen2024longvu, liang2024keyvideollmlargescalevideokeyframe], rely on vision–text similarity models like CLIP for keyframe selection. While efficient, these methods perform static, one-shot selection, limiting their adaptability to complex or evolving queries. Agentic approaches [wang2025videochata1thinkinglongvideos, liu2025videomindchainofloraagentlong] address this limitation through iterative frameworks that dynamically refine frame selection. For example, VideoAgent [wang2024videoagentlongformvideounderstanding] employs an LLM as a central agent in a state–action–observation loop, whereas VideoTree [wang2025videotreeadaptivetreebasedvideo] adopts a hierarchical tree structure to perform coarse-to-fine search, directing computation toward the most relevant segments. More recent model-internalized selection methods [tong2025thinkingvideovideogeneration] embed frame selection directly into the model’s reasoning process. GenS [yao2025generativeframesamplerlong] trains a lightweight generative sampler to identify query-relevant frames end to end, though the selection and QA stages remain decoupled.
Our proposed REVISOR framework, after reinforcement learning training, can autonomously identify and explore visual content that requires additional careful review during the inference stage. Unlike previous methods that are limited to statically searching for important video segments based solely on the question itself, our approach fully integrates with the reasoning capabilities of MLLMs. This allows it to engage in deeper deliberation and, through comprehensive interaction with its ongoing reasoning, more precisely pinpoint critical visual information, thereby significantly enhancing the model’s reasoning ability.
Self-Reflection Mechanism. Incorporating self-reflection into inference has been shown to improve model performance [guo2025deepseek]. Self-Refine [madaan2023self] implements an iterative feedback–revision loop without requiring additional training or external supervision. SCoRe [kumar2024training] adopts multi-round reinforcement learning to cultivate reflective reasoning skills, yielding substantial gains. GEPA [agrawal2025gepa] further introduces a Pareto-based reflective prompt-evolution framework that outperforms reinforcement learning in both sample efficiency and overall performance.
Extending self-reflection to multimodal settings likewise yields consistent improvements [yang2025look, jian2025look]. VL-Rethinker [wang2025vl] incorporates a self-verification–and-correction stage via reinforcement learning, substantially enhancing multimodal reasoning. SRPO [wan2025srpo] further strengthens self-reflection and self-correction by curating reflection-oriented supervised data and introducing a reflection-aware reward under group relative policy optimization, significantly boosting complex multimodal reasoning.
However, existing multimodal self-reflection mechanisms still depend on text-based re-evaluation processes. They lack explicit reflective operations over visual information and therefore remain fundamentally text-centric. The REVISOR framework proposed in this paper addresses this limitation by extending text-centric reflection into a truly multimodal reflective process.
@@ 7 Conclusion
[…]
Supplementary Material
8 More Experimental Details on REVISOR
In Sec. 8.1, we present the detailed experimental setup for experiments involving the REVISOR framework. Sec. 8.2 outlines the composition of the training data for the REVISOR framework. Sec. 8.3 then presents supplementary experimental results, including comprehensive results on the Temporal Video Grounding task and complete ablation studies on the reward scaling factors $\lambda_{1}$ and $\lambda_{2}$.
Table 5: Comprehensive evaluation of REVISOR on the temporal video grounding task. Bold text indicates the best performance.
Charades-STA NExT-GQA
Model R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 mIoU
Qwen2.5-VL-7B[bai2025qwen2] 57.1 33.6 15.5 36.9 31.6 18.1 7.5 20.9
VTimeLLM[huang2024vtimellm] 55.3 34.3 14.7 34.6 37.9 20.2 9.7 24.4
iMOVE[li2025imove] 71.7 51.3 26.1 47.3 - - - -
TimeChat[ren2024timechat] 51.5 32.2 13.4 - 34.1 17.9 6.2 20.6
VideoChat-TPO[yan2025task] 58.3 40.2 18.4 38.2 41.2 23.4 8.2 27.7
TVG-R1[chen2025datasets] 70.8 50.5 23.9 46.7 41.7 20.8 10.0 29.3
Ours 76.5 57.3 31.8 51.4 47.6 25.5 11.9 33.2
Table 6: Complete ablation study of the Dual Attribution Decoupled Reward mechanism. Bold fonts highlight the best performance. The row marked in gray represents our base model.
$\boldsymbol{\lambda_{1}(R_{final})}$ $\boldsymbol{\lambda_{2}(R_{causal})}$ VideoMME LongVideoBench LVBench MLVU NExT-GQA
- - 64.3 56.5 40.2 67.3 20.87
0.3 0.6 64.0 56.0 41.1 68.7 33.9
0.6 0.0 62.2 54.0 40.8 68.3 32.1
0.45 0.45 64.3 57.3 41.2 68.5 33.7
0.5 0.4 64.9 57.1 41.8 69.0 33.4
0.6 0.3 65.7 57.5 42.0 69.8 33.2
0.8 0.1 64.6 57.6 40.3 68.4 31.6
Figure 6: REVISOR framework training dataset composition. The training dataset for the REVISOR framework consists of three tasks: Short Video QA, Temporal Grounding, and Grounded Video QA, totaling 25K training samples. The parenthetical value for each dataset denotes its specific sample contribution.
8.1 Detailed Experimental Setup
Our experiments utilized Qwen2.5-VL-7B as the base model, which comprises a visual encoder, a merger projector, and a large language model. REVISOR was trained using a single stage of reinforcement learning, eliminating the need for an additional supervised fine-tuning phase. We extended the verl framework to support REVISOR’s training. Following the approach in DAPO, we removed the KL regularization term from GRPO.
Key hyperparameters were set as follows: $\lambda_{1}=0.6$, $\lambda_{2}=0.3$, a learning rate of $1\times 10^{-6}$, a batch size of 32, and 8 rollouts. During both training and evaluation, input video tokens were limited to a maximum of 8192, sampled at 1 FPS. For the Review Segment, the sampling rate was increased to 2 FPS, while still adhering to a maximum of 8192 video tokens. Absolute timestamps were displayed in the lower-left corner of each image frame. The entire training phase consisted of 792 optimization steps.
8.2 Detailed Information of Training Data
As detailed in Fig. 6, our training corpus spans three tasks central to video understanding: Short Video QA, Temporal Grounding, and Grounded Video QA. For Short Video QA, we aggregate data from LLaVA-Video-178K, CLEVRER, NExT-QA, PerceptionTest, and STAR, all of which provide short video clips ($\leq$ a few minutes). From these datasets, we sample a balanced set of 20K video–question–answer triplets. The Temporal Grounding task is sourced directly from TimeRFT without further filtering. For Grounded Video QA, we incorporate CG-Bench and ReXTime, from which we randomly select 2K examples.
8.3 Supplementary Experimental Results
Temporal Video Grounding. As summarized in Tab. 5, we conduct an extensive assessment of REVISOR’s temporal grounding performance across both the Charades-STA and NExT-GQA benchmarks. On Charades-STA, REVISOR attains an R@0.3 accuracy of 76.5%, surpassing the prior state-of-the-art method iMOVE by a margin of 4.8%. On NExT-GQA, REVISOR reaches an R@0.5 accuracy of 25.5%, outperforming TVG-R1, designed explicitly with reinforcement learning to enhance temporal localization, by 4.3%. Together, these results highlight the robustness and effectiveness of REVISOR, equipped with DADR, in achieving precise temporal video localization.
Ablation Study of DADR Mechanism. We conducted an in-depth analysis of the Dual Attribution Decoupled Reward (DADR) mechanism. As shown in Tab. 6, when $\lambda_{2}=0$, that is, when only the Final Answer Verification Reward is applied, the performance of REVISOR on Video-MME decreases substantially from 65.7% to 62.2%, even falling below the base model. This demonstrates that, without the Causal Segment Sufficiency Reward (CSSR), the model struggles to reliably identify the correct review segment $S$ due to the sparsity of the reward signal.
When $\lambda_{2}$ exceeds $\lambda_{1}$, the model’s temporal grounding improves; however, its long-video reasoning capability degrades. For example, MLVU performance declines from 69.8% to 68.7%, suggesting that the model overemphasizes locating the review segment $S$ while underutilizing it for answer derivation. Consequently, we set $\lambda_{1}>\lambda_{2}$ to encourage accurate localization of $S$ while still promoting strong reasoning based on it.
Nonetheless, if $\lambda_{1}$ is too large and $\lambda_{2}$ too small, the model’s ability to locate $S$ deteriorates, harming long-form video understanding. Conversely, when $\lambda_{1}$ is only slightly greater than $\lambda_{2}$, REVISOR identifies $S$ effectively but fails to fully leverage it for reasoning, resulting in strong temporal grounding but weaker long-video comprehension. Empirically, we find that $\lambda_{1}=0.6$ and $\lambda_{2}=0.3$ provide an effective balance, enabling REVISOR to both accurately localize $S$ and utilize it to enhance reasoning performance.
9 Training Qwen2.5-VL-7B with Textual Reflection Mechanism on Video Data
To ensure a fair comparison, we train a Qwen2.5-VL-7B model equipped with a text-based self-reflection mechanism using the datasets listed in Fig. 6. Specifically, unlike REVISOR, the text-reflection model generates only textual output during the reflection phase. Apart from this distinction, all other experimental settings remain identical to those of REVISOR. The learning rate is set to $1\times 10^{-6}$, the total batch size is 32, and the number of rollouts is set to 8. In addition, the total number of input video tokens is capped at 8192. The prompt template is shown in Fig. 11.
10 Case Study of REVISOR Framework
Improved Video Reasoning Capability. REVISOR can significantly enhance the long-form video understanding capabilities of MLLMs. Fig. 7, Fig. 8, and Fig. 9 respectively demonstrate these improvements from three perspectives: more precise detail capture, more comprehensive scene understanding, and more accurate object counting.
Visual Reflection is Even More Important. In this paper, we repeatedly emphasize that, for long-video understanding tasks, visual reflection is more important than textual reflection. In Sec. 5.1, we quantitatively validate this conclusion by monitoring changes in the length of textual reflections during training. Here, we further illustrate this point with a concrete example. As shown in Fig. 10, during training, the textual reflection generated by the REVISOR framework for a given question becomes increasingly concise, while its retrieval of key video segments becomes increasingly accurate. Ultimately, the model precisely identifies the critical 25-second segment within a 300-second video, confirming that respiratory cells could potentially be used in treatment for cardiovascular disease.
Figure 7: Successful example of the REVISOR framework: achieving more accurate detail capture.
Figure 8: Successful example of the REVISOR framework: achieving more accurate scene understanding.
Figure 9: Successful example of the REVISOR framework: achieving more accurate object counting.
Figure 10: As training progressed, REVISOR’s outputted reflective text paragraphs became increasingly concise, and the video segments it referenced grew more precise. Ultimately, it accurately concluded that respiratory cells could potentially be used in the treatment of cardiovascular disease.
11 Prompt Templates of REVISOR
The complete prompt template used during the training of the REVISOR framework consists of three primary components: the system prompt, the initial reasoning stage, and the reflective reasoning stage. Fig. 12 illustrates the templates for both the initial reasoning and reflective reasoning stages, while Fig. 13 presents the system prompt template. A key distinction from the plain-text reflection mechanism is that, during the initial reasoning stage, the REVISOR framework identifies the critical video segments that require further examination.
The Complete Prompt Template for Plain-Text Reflection Mechanism System Prompt:
You are a helpful assistant.
Stage 1: Preliminary Reasoning Phase
User Instruction:
<User Question>. Please think carefully and then provide your answer.
Output Format:
Format strictly as: <think> Your reasoning steps </think> <answer> Your answer </answer>
Stage 2: Reflective Reasoning Phase
User Instruction:
Please rethink the reasoning process above and provide your final answer.
Output Format:
Format strictly as: <rethink> Your rethinking reasoning steps </rethink> <answer> Your final answer </answer>
Figure 11: The complete prompt template used during the training of Qwen2.5-VL-7B equipped with a pure text-based reflection mechanism, covering both the initial reasoning and reflective reasoning stages.
Reasoning Prompt Template for REVISOR Framework Training Stage 1: Preliminary Reasoning Phase
User Instruction:
<User Question>. Please think first, and then use the temporal_zoom_in_tool to find the video segment that can answer the user’s question.
Output Format:
Format strictly as: <think> Your reasoning steps </think> <time_interval>[start_time, end_time] </time_interval>
Stage 2: Reflective Reasoning Phase
User Instruction:
Please refer to the Visual Review segment above, think carefully, and provide your final answer.
Output Format:
Format strictly as: <rethink> Your reasoning steps </rethink> <answer> Your final answer </answer>
Figure 12: Reasoning prompt template for REVISOR framework training
System Prompt Template for REVISOR Framework Training You are a helpful assistant. # Tools
You may call one or more functions to assist with the user query. You are provided with function signatures within <tools></tools> XML tags:
<tools> 1{ 2 "type": "function", 3 "function": { 4 "name": "temporal_zoom_in_tool", 5 "description": "Identify the precise time segment in the video that contains enough information to answer the question.", 6 "parameters": { 7 "properties": { 8 "interval": { 9 "type": "str", 10 "description": "The time range to zoom in on, formatted as start_time to end_time’. Timestamps are in seconds." 11 } 12 }, 13 "required": ["interval"] 14 } 15 } 16} </tools>
# Explanation of ”temporal zoom-in”
When you call temporal_zoom_in_tool with <time_interval> [start_time, end_time] </time_interval>, the tool returns a new sequence of denser video frames sampled from the specified time range ([start_time, end_time]) in the original video. This provides higher temporal precision, helping you answer the user’s question more accurately.
# How to call a tool
Return the interval directly in XML tags: <time_interval> [12.3, 28.7] </time_interval>
Figure 13: System prompt template for training the REVISOR framework.